The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #79

April 23, 1999 - 1:00 p.m. EST (10:00 a.m. PST)


The Lunar Prospector spacecraft continues to collect good science data in the extended mission orbit. All five science instruments are on: the Gamma Ray Spectrometer (GRS), the Neutron Spectrometer (NS), the Magnetometer (MAG), the Electron Reflectometer (ER), and the Alpha Particle Spectrometer (APS). Data continues to be collected for the Doppler Gravity Experiment (DGE) as well.

Investigation into the commanding anomaly continues. Since the last report, only one command (out of a total of 72) has not executed normally. Again, this command did not fire when the downlink signal was showing lots of noise and the spacecraft was in the northern hemisphere. It is believed that the uplink signal may occasionally be corrupted because of coherent reflections from the surface of the moon.

During the last week, there have been 4 (of 11) occurrences of the battery voltage reaching the yellow low-voltage limit during an eclipse while the propellant tank heater was on.

The extended mission orbit maintenance burn was successful executed on April 20 (DOY 111 GMT). There were no command anomalies during this event. A timeline is given below.

111/01:00 Turn Earth-Moon Sensor (EMS) ON
111/01:01 Turn Pressure Transducer ON
111/01:06 Turn A1/A2 heaters ON
111/01:08 Load maneuver parameters
111/01:32 Fire delta-V part 1 (A1/A2 for 45.2 sec, 7.10 m/s burn, 0.57 kg fuel)
111/01:36 Turn A1/A2 heaters ON
111/01:43 Load maneuver parameters
111/02:14 Fire delta-V part 2 (A1/A2 for 45.3 sec, 7.12 m/s burn, 0.56 kg fuel)
111/02:17 Turn A1/A4 heaters ON
111/02:53 Load maneuver parameters
111/02:58 Fire reorientation (A1/A4 for 4 pulses, 0.7 deg precession, 0.01 kg)
111/03:01 Turn T1 heater ON
111/03:16 Load maneuver parameters
111/03:19 Fire spin trim (T1 for 0.63 sec, 0.155 rpm change, 0.003 kg fuel)
111/03:21 Turn Pressure Transducer OFF

The command failure occurred when the EMS was being turned off the day after the maneuver. Resending the command turned the sensor off.

On April 21, the Alpha Particle Spectrometer was turned on. It had been turned off since February 8 when noise from this instrument was seen to interfere with the Neutron Spectrometer. The APS is currently collecting good data and does not seem to be generating noise. The data will continue to be monitored.



Current spacecraft state (0000 GMT 4/23/99, DOY 113):

Orbit: 5773
Downlink: 3600 bps
Spin Rate: 12.00 rpm

Spin Axis Attitude (ecliptic):
Latitude: -88.5 deg
Longitude: 040 deg

Trajectory:
Periapsis Alt: 21.6 km
Apoapsis Alt: 38.4 km
Period: 111 min

Occultations: 39 minutes
Eclipses: 49 minutes
Propellant Remaining: 12.47 kg

The next orbit maintenance burn is scheduled for May 5. This maneuver will put the spacecraft into a similar 30-km average altitude orbit, but one that will have periapsis on the far side of the moon for most of the time, rather than the near side of the moon as the current orbit has. Maneuvers will then continue once every 28 days through the end of mission.